### v1

# Understanding and Mitigating Tokenization Bias in Language Models

Abstract
State-of-the-art language models are autoregressive and operate on subword units known as tokens. Specifically, one must encode the conditioning string into a list of tokens before passing to the language models for next-token prediction. We show that, for encoding schemes such as maximum prefix matching, tokenization induces a sampling bias that cannot be mitigated with more training or data. To counter this universal problem, we propose a novel algorithm to obtain unbiased estimates from a model that was trained on tokenized data. Our method does not require finetuning the model, and its complexity, defined as the number of model runs, scales linearly with the sequence length. As a consequence, we show that one can simulate token-free behavior from a tokenized language model. We empirically verify the correctness of our method through a Markov-chain setup, where it accurately recovers the transition probabilities, as opposed to the conventional method of directly prompting tokens into the language model.
Machine Learning, ICML
[…]
1 Introduction
Tokenization is a preprocessing procedure used in many state-of-the-art (SOTA) language models (LMs) such as GPTs (Brown et al., 2020), Llama (Touvron et al., 2023) and Gemini (Gemini, 2023). It divides the input text into smaller subword units while retaining linguistic importance, helping to address vocabulary limitations such as unknown words. Tokenization also shortens (compresses) the input context length (Sennrich et al., 2015; Kudo & Richardson, 2018). Since effective compression allows transformer-based LMs to handle longer context strings, many works (Zouhar et al., 2023; Gallé, 2019; Goldman et al., 2024) have focused on enhancing vocabulary design and encoding algorithms for better performance in downstream tasks. However, the relationship between compression and model performance remains unclear. Some research suggests the impact of compression is not always positive (Schmidt et al., 2024; Dagan et al., 2024; Goyal et al., 2023). Consequently, understanding tokenization’s effect on model performance continues to be an open question.
Tokenization has been criticized for introducing many shortcomings in LMs. These include sensitivity to spelling and morphological structure (Xue et al., 2022), language-based biases (Petrov et al., 2024), subpar performance in specific tasks such as arithmetic (Singh & Strouse, 2024), or new domains (Liu et al., 2023a). One approach to address these issues is through fine-tuning the model with new vocabularies; however, this often complicates the training process and requires domain-specific expertise (Chen et al., 2023; Liu et al., 2023b). Furthermore, the performance gains do not provide a theoretical understanding of whether these limitations truly arise from the tokenization process or result from suboptimal model training. Another direction is to develop token-free LMs (Yu et al., 2024; Nawrot et al., 2022; Tay et al., 2021). While this approach has potential as it eliminates tokenization-related issues, it significantly increases the context length, resulting in performance that still lags behind the SOTA tokenized LMs  [^1] We refer language models that process tokenized texts as tokenized language models (tokenized LMs).(Yu et al., 2024).
In this work we offer new theoretical insights on the behavior of tokenized LMs. We show that they are statistically equivalent to their token-free counterparts. Specifically, we examine the maximum prefix encoding scheme employed in the WordPiece tokenization method (Devlin et al., 2018; Song et al., 2020) and find that this process not only results in biased estimates of next token probabilities, but also leads to overall skewed estimates of subsequent character probabilities. In general, this bias persists despite an increase in training data, even within the simple setting of a 1st-order Markov chain. Such bias occurs due to the implicit disparity between the domain of the conditioning context, namely, characters versus tokens. Nevertheless, we will show that it is possible to correct this bias without resorting to finetuning. Once adjusted, it becomes possible to simulate the token-free behavior learned implicitly by the tokenized LM and even (theoretically) mimic the behavior of another tokenized model employing a distinct vocabulary set, all without requiring finetuning. Our specific contributions are as follows:
  • We show the presence of a bias in the next-token distribution that arises as a result of the tokenization process.
  • We present a novel algorithm capable of transforming a tokenized LM into its token-free counterpart, thereby effectively eliminating the sampling bias.
  • We verify the correctness of our algorithm on learning the transition matrix of a $k$-th order Markov chain.
2 Problem Setup
[…]
ID Token
1 $A$
2 $B$
3 $AA$
[…]
$AA$$A$$B$Output Tokens: ``${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}AA}|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}B}|{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}A}|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}B}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}AA}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}AA}|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}B}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}AA}|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}B}"$$\alpha\beta$1.01.01.01.00.00.00.00.00.00.00.00.0${(}1{-}\alpha{)}^{2}$$(1{-}\alpha)\beta$$\alpha{(}1{-}\alpha{)}$$\alpha$$1{-}\beta$
Figure 1: Next-Character sampling bias introduced by the WordPiece encoding algorithm. In this example, given the context token ``$A"$, the model will always predict the next token as ``$B"$ with probability $1.0$. We present a technique that, given a language model trained on tokenized domain, eliminate this bias and recover the accurate unbiased sampling distribution.
We begin by establishing the tokenization and language models setup in our paper. We then describe the next-character sampling bias problem due to tokenization.
2.1 Notations and Setup.
String Notations. For any string $s$, we denote its substring from $i$ to $j$ as $x^{j}_{i}{\vcentcolon=}{x_{i}x_{i+1}..x_{j}}$, where each $x$ is a character of the alphabet $\mathcal{A}$. For a given string $x^{N}_{1}$, we define the prefix function that generates a set containing all possible prefix strings of $x^{N}_{1}$, represented as $\mathrm{prefix}(x^{N}_{1}){=}\{x^{1}_{1},x^{2}_{1},x^{3}_{1},...,x^{N}_{1}\}$. Also, we define a concatenation function $\mathrm{concat}(.)$ that concatenates the given list of strings, e.g given $s_{1}{=}x^{N_{1}}_{1}$ and $s_{2}{=}y^{N_{2}}_{1}$, we obtain $\mathrm{concat}(s_{1},s_{2}){=}\mathrm{concat}(x^{N_{1}}_{1},y^{N_{2}}_{1}){=}x_{1}...x_{N_{1}}y_{1}...y_{N_{2}}$. Finally, we denote the set of all strings that start with a prefix $x^{n}_{1}$ as $\mathcal{S}(x^{n}_{1}){=}\{s|x^{n}_{1}{\in}\mathrm{prefix}(s)\}$.
Tokenization Setting. We assume having a predefined vocabulary $\mathcal{V}$ constructed using any tokenization algorithm such as Byte-Pair-Encoding (BPE) or Lempel-Ziv-Welch (LZW), with the condition that $\mathcal{A}\subseteq\mathcal{V}$. We use $t$ to denote a token in $\mathcal{V}$, i.e. $t\in\mathcal{V}$. Importantly, we use the longest prefix matching strategy for tokenization (encoding), denoted as $\mathrm{encode}(.)$, similar to the approach used in the Wordpiece algorithm (Devlin et al., 2018; Song et al., 2020). Given a sequence of tokens $t^{k}_{1}$, the function $\mathrm{decode}(.)$ returns the concatenated string resulting from processing each token in the sequence. Finally, the set of all strings that starts with the tokens $t^{k}_{1}$ is defined as $\mathcal{S}(t^{k}_{1}){=}\{s|t^{k}_{1}{=}\mathrm{encode}(s)^{k}_{1}\}$.
Tokenized LMs. We assume having access to a tokenized autoregressive LM with parameters $\theta$ that is trained with tokens from $\mathcal{V}$ and maximum prefix matching. The target distributions on the character domain is denoted as $P_{\mathrm{gt}}(x^{N}_{n+1}|x^{n}_{1})$ and on the token domain is $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$. For simplicity, unless otherwise stated, we implicitly assume each probability term involves $\theta$. Using the model, we assume that one can compute $P(t_{i+1}|t^{i}_{1})$ for any integer $i>0$. In this work, we consider LMs trained under the standard setup, where each string $s$ in the dataset is first tokenized with the encoding function $\mathrm{encode}(.)$ and vocabulary $\mathcal{V}$, and the parameters $\theta$ are optimized to maximize the predictive likelihood of the next token in the tokenized dataset.
2.2 Next-Character Sampling Bias
We first define the (next-character) sampling bias problem that describes the discrepancy between the character level and token level predictions for tokenized LMs.
Definition 2.1.
(Next-Character Sampling Bias) Let the input prompt string $x^{n}_{1}$ has $t^{i}_{1}{=}\mathrm{encode}(x^{n}_{1})$ as the corresponding encoding. The next-character sampling bias occurs for this prompt when $P_{\mathrm{gt}}(x_{n+1}|x^{n}_{1}){\neq}P_{\mathrm{gt}}(x_{n+1}|t^{i}_{1})$ where $P_{\mathrm{gt}}(x_{n+1}|t^{i}_{1}){=}\sum_{t\in\mathcal{E}}P_{\mathrm{gt}}(t_{i+1}{=}t|t^{i}_{1})$ where =$\mathcal{E}=\{t\in\mathcal{V}|\mathrm{decode}(t)\in\mathcal{S}(x_{n+1})\}$.
In other words, the probability of the next character being “c” may be different from the sum of the probabilities of all tokens that start with “c”. Note that this character-level probability offers a broader perspective compared to the probability of the subsequent token being exactly “c”.
Example. Consider a first order Markov chain with two states {``$A",``B"\}$ as shown in Figure 1 (left). Each string is tokenized with $\mathcal{V}{=}\{``AA",``A",``B"\}$, which leads to a new Markov chain whose states and transition matrix is shown in Figure 1 (right). Details on computing the transition matrix of the new Markov chain is in Appendix F. We first observe that for the prompt $s_{1}{=}``AA"$ and $s_{2}{=}"B"$, there is no bias problem after marginalization [^2] For example, we have $P_{\mathrm{gt}}(t_{i{+}1}{=}``AA"|t_{i}{=}``AA")+P_{\mathrm{gt}}(t_{i{+}1}{=}``A"|t_{i}{=}``AA")=\alpha=P_{\mathrm{gt}}(x_{n{+}1}{=}``A"|x_{n}{=}``A")$. However, for the prompt $s_{3}{=}``A"$, the sampling bias occurs as $P_{\mathrm{gt}}(x_{2}{=}``B"|t_{1}{=}``A"){=}1.0$, which is not equal to $P_{\mathrm{gt}}(x_{2}{=}``B"|x_{1}{=}``A"){=}\alpha$, i.e. the optimally trained LM will always output ``$B"$. In fact, for any context string that ends with token ``$A"$, e.g ``$AA|A"$ and ``$B|A"$ (tokens are separated by ``|"conditional``"``|"` ` | "), such LM will always output ``$B"$.
Since this applies to any optimally trained LM, increasing the training set size does not mitigate this problem. The reason for this sampling bias is that, during the tokenization process with longest prefix matching, the token ``$A"$ must be followed by the token ``$B"$. Otherwise, the encoding process will merge to create a longer token ``$AA"$. To generalize this phenomenon, we start with the definition of invalid encodings.
Definition 2.2.
(Invalid Encodings) The list of tokens (an encoding) $t^{i}_{1}$ is invalid if $\mathrm{encode}(\mathrm{decode}(t^{i}_{1})){\neq}t^{i}_{1}$. Otherwise, it is a valid encoding.
For example, let $\mathcal{V}{=}\{``c",``a",``t",``at",``cat"\}$ then [``$c"{,}``at"{,}``t"]$ and [``$c"{,}``a"{,}``t"{,}``t"]$ are invalid encodings of ``$catt"$. We now show in Proposition 2.3 that the existence of invalid encodings introduces sampling bias, generalizing the observed phenomenon in the Markov chain example to any autoregressive distribution.
Proposition 2.3.
(Token-Induced Zero Probability) Let $t^{i}_{1}$ be a sequence of input tokens. For any invalid encoding $t^{i}_{1}$, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ and the conditional probability $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ is undefined. In the case $t^{i}_{1}$ is valid, then $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1}){=}0.0$ if $t^{i+1}_{1}$ is invalid. Furthermore, let $x^{n}_{1}{=}\mathrm{decode}(t^{i}_{1})$, then for any string $x^{N}_{n+1}$ such that $\mathrm{encode}(\mathrm{concat}(\mathrm{decode}(t^{i}_{1}),x^{N}_{n+1})){\neq}t^{i}_{1}$, then $P_{\mathrm{gt}}(x^{N}_{n+1}|t^{i}_{1}){=}0.0.$
Proof.
See Appendix C. ∎
Remark 1.
Proposition 2.3 implies that LMs may not function as expected when presented with invalid encodings, because these models will never be exposed to such inputs within the dataset. This directly implies that the practice of evaluating LMs under different encodings (Cao & Rimell, 2021; Chirkova et al., 2023) is suboptimal.
3 Alleviating Sampling Bias
We propose a method to remove the described bias and recover the original token-free autoregressive model, i.e. expressing the implicitly learned $P(x^{N}_{n+1}|x^{n}_{1})$ using the tokenized LM that outputs the conditional probability $P(t_{i+1}|t^{i}_{1})$. For $N{=}n{+}1$, this captures the behavior of a token-free model, i.e. sampling the next character instead of a whole token. We assume our LM follows Proposition 2.3 on zero probability events and undefined conditional probability for invalid encodings. Appendix G justifies this assumption and provides its practical implementation.
Our method consists of two stages. In the first stage, the idea is to identify the condition when $P(x^{N}_{n+1}|t^{i}_{1})=P(x^{N}_{n+1}|x^{n}_{1})$ where $t^{i}_{1}=\mathrm{encode}(x^{n}_{1})$. Once identified, we can refactor the conditional probability to match the conditioning events. In the second stage, we compute $P(x^{N}_{n+1}|t^{i}_{1})$ using the LM output probability, i.e. $P(t_{i+1}|t^{i}_{1})$, through the novel Branch and Pass (BPTree) Algorithm.
3.1 Refactoring
Our method removes the bias by connecting character and token domains through a special subset of tokens $\mathcal{V}^{*}{\subset}\mathcal{V}$, whose elements $t^{*}{\in}\mathcal{V}^{*}$ are not a substring of any other tokens in $\mathcal{V}$ but itself. For example, given $\mathcal{V}{=}\{``AAA",``AA",``CB",``A",``B",``C"\}$, then $\mathcal{V}^{*}=\{``AAA",``CB"\}$. In the Markov example in Section 2, this corresponds to the tokens ``$AA"$ and ``$B"$. Also, we assume that any string $x^{N}_{1}$ has the first token $t_{1}\in\mathcal{V}^{*}$ [^3] This is not restrictive, as many current language models starts with a special beginning token in $\mathcal{V}^{*}$, e.g. the start token ${<}start{>}$ in SentencePiece (Kudo & Richardson, 2018).. Consider the input string $x^{n}_{1}$ and its corresponding encoding $t^{i}_{1}{=}\mathrm{encode}(x^{n}_{1})$, Proposition 3.1 shows the sufficient condition for $\mathcal{S}(t^{i}_{1}){=}\mathcal{S}(x^{n}_{1})$.
Proposition 3.1.
Let $s^{*}=x^{n}_{1}$, where $t^{i}_{1}=\mathrm{encode}(s^{*})=\mathrm{encode}(x^{n}_{1})$. Then we have $\mathcal{S}(t^{i}_{1})\subset\mathcal{S}(x^{n}_{1})$, i.e. for any string $s$ where $t^{i}_{1}=\mathrm{encode}(s)^{i}_{1}$, we have $P(x^{n}_{1}|t^{i}_{1})=1.0$. In the case $t_{i}\in\mathcal{V}^{*}$, then we also have that $\mathcal{S}(t^{i}_{1})=\mathcal{S}(x^{n}_{1})$, i.e. any string $s$ where $x^{n}_{1}\in\mathrm{prefix}(s)$ must have the first $i$ tokens as $t^{i}_{1}$ and $P(t^{i}_{1}|x^{n}_{1})=1.0$.
Proof.
See Appendix D. ∎
The intuition for Proposition 3.1 is that the subsequent string after $t_{i}{\in}\mathcal{V}^{*}$ cannot change the tokenization for $x^{n}_{1}$. We now establish one of the main results in Corollary 3.2.
Corollary 3.2.
Following Proposition 3.1, suppose $t_{i}\in\mathcal{V}^{*}$ then we have $P(x^{N}_{n+1}|x^{n}_{1}{)}{=}P(x^{N}_{n{+}1}|t^{i}_{1})$. Similarly, we also have $P(t^{j}_{i+1}|x^{n}_{1}{)}{=}P(t^{j}_{i+1}|t^{i}_{1})$.
Proof.
See Appendix D. ∎
We note that Proposition 3.1 and Corollary 3.2 always hold, regardless of the value of $\theta$. In general, consider when the last token of $\mathrm{encode}(x^{n}_{1})$ is not in $\mathcal{V}^{*}$, we can refactor $P(x^{N}_{n+1}|x^{n}_{1})$ as follow:
$P(x^{N}_{n+1}|x^{n}_{1}){=}\frac{P(x^{N}_{n_{k}+1}|t^{k}_{1})}{P(x^{n}_{n_{k}+1}|t^{k}_{1})},$ (1)
where $k$ is the last token in $\mathrm{encode}(x^{n}_{1})$ such that $t_{k}{\in}\mathcal{V}^{*}$ and $x^{n_{k}}_{1}{=}\mathrm{decode}(t^{k}_{1})$, where $n_{k}\leq n$. Proof details of this step can be found in the Appendix E. We then use the BPTree algorithm to compute each term in the RHS individually.
3.2 The Branch and Pass Algorithm
Algorithm 1 Branch and Pass Algorithm on Discrete Probability Tree. This algorithm recursively computes $P(x^{N}_{n_{k}+1}|t^{k}_{1})$.
1:procedure BPTree($x^{N}_{n_{k}+1},t^{k}_{1}$)
2:    // Branching Step:
3:    =$\mathcal{B}=\{t\in\mathcal{V}|x^{N}_{n_{k}{+}1}{\in}\mathrm{prefix}(\mathrm{decode}({t}))\}$
4:    $\mathrm{b_{val}}={\sum\limits_{t{\in}\mathcal{B}}}P(t_{k{+}1}=t\big{|}t^{k}_{1})$
5:    // Base Case:
6:    if $\mathrm{encode}(x^{N}_{i})\in\mathcal{V}$ then
7:        return $\mathrm{b_{val}}$
8:    end if
9:    //Extract the Next Token:
10:    $t_{k+1}=\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$
11:    // Passing Step:
12:    $\mathrm{p_{val}}=P(t_{k+1}\big{|}t^{k}_{1})$
13:    $\mathrm{p_{val}}=\mathrm{p_{val}}\times\text{BPTree}(x^{N}_{n_{k+1}+1},t^{k+1}_{1})$
14:    return $\mathrm{b_{val}}+\mathrm{p_{val}}$
15:end procedure
We present the BPTree algorithm in Algorithm 1, that allows us to compute the probabilities $P(x^{N}_{n_{k}+1}|t^{k}_{1})$ and $P(x^{n}_{n_{k}+1}|t^{k}_{1})$ in Equation (1). Note that this algorithm does not require $t_{k}{\in}\mathcal{V}^{*}$. Details on the algorithmic correctness are shown in Appendix E.
The idea is to marginalize out $P(x^{N}_{n_{k}+1}|t^{k}_{1})$ by considering two complementary events: when the next token $t_{k+1}$ has a prefix $x^{N}_{n_{k}+1}$ ($\mathrm{b_{val}}$ in the Branch Step) versus when the next token $t_{k+1}$ is contained within $x^{N}_{n_{k}+1}$ ($\mathrm{p_{val}}$ in the Pass Step). Formally, BPTree computes the following probabilities:
$\vspace*{-20pt}\mathrm{b_{val}}$ $=P(x^{N}_{n_{k}{+}1},t_{k+1}\in\mathcal{B}(x^{N}_{n_{k}+1}))\big{|}t^{k}_{1}),$ (2)
$\mathrm{p_{val}}$ $=P(x^{N}_{n_{k}{+}1},t_{k+1}\notin\mathcal{B}(x^{N}_{n_{k}+1}))\big{|}t^{k}_{1}),\vspace{-20pt}$ (3)
where $\mathcal{B}(x^{N}_{n_{k}+1}){=}\{t{\in}\mathcal{V}|x^{N}_{n_{k}{+}1}{\in}\mathrm{prefix}(\mathrm{decode}({t}))\}$ and we immediately see that $P(x^{N}_{n_{k}+1}|t^{k}_{1}){=}\mathrm{b_{val}{+}p_{val}}$.
We provide an intuitive explanation for the algorithm following the example in Figure 2. Here, we would like to compute the probability $P(x^{n_{k}+3}_{n_{k}+1}{=}``bee"|t^{k}_{1})$. The first possibility is that ``$bee"$ is a prefix of the next token, so we search for all such tokens (line 3 in the algorithm) and sum up their probability (line 4), i.e. $\mathrm{b_{val}}{=}P(t_{k+1}{=}``beer"|t^{k}_{1})$. Figure 2 visualizes this step as branching out the tree by finding all tokens completing the string. Since ``$beer"$ is not the only string that contains ``$bee"$, e.g. ``$beep",``been"$, etc. we need to compute the probability for these other scenarios, each of which has $t_{k+1}{=}``b"$ (the first token in ``$bee"$, line 10 and 12) due to maximum prefix encoding. Then, we want to compute the probability that the subsequent string is ``$ee"$ (line 13), given the previous $t^{k}_{1}$ and $t_{k+1}{=}``b"$, which is the output of the BPTree algorithm but for $x^{n_{k}+3}_{n_{k}+2}{=}``ee"$ and $t^{k+1}_{1}$. Formally, in the Passing step: $\mathrm{p_{val}}{=}P(t_{k+1}{=}``b"|t^{k}_{1})P(x^{n_{k}+3}_{n_{k}+2}{=}``ee"|t^{k}_{1},t_{k+1}{=}``b")$. We continue the procedure until meeting the base case, where the string must be a prefix of the next token (usually, when there is only a single character left). Finally, by computing the sum of the branch and pass steps, we obtain the desired conditional probability $\mathrm{b_{val}}{+}\mathrm{p_{val}}{=}P(x^{n_{k}+3}_{n_{k}+1}{=}``bee"|t^{k}_{1})$.
4 Experiments
Figure 2: BPTree Visualization. At each recursive call, the Branch step finds tokens that starts with the query string while the Pass step extracts and employs the next token and leftover string for the next recursive call until meeting the base case.
[…]
Figure 3: Our method accurately estimates the transition probability of a 3rd order Markov chain while the baseline method fails to.
5 Conclusion
This work identifies the next-character sampling gap between a tokenized model and a token-free one, which persists even for optimally trained models. We present a probabilistic approach to effectively eliminate this bias without requiring additional training. This closes the sampling gap between tokenized and token-free models, suggesting that language models implicitly absorb character-level information despite being trained solely on tokenized text. This result implies that it is theoretically possible to simulate the behavior of another language model trained using different vocabulary without any fine-tuning, since it is possible to transfer from token-free models to tokenized counterparts. Future work seeks to extend our investigation for alternative encoding strategies, most notably the widely used BPE tokenization (Sennrich et al., 2015), in state-of-the-art models.
Appendix A Related Work
Theory of Tokenization. Existing works on tokenization generally support the idea that compressing tokens enhances model performance (Gallé, 2019; Gutierrez-Vasques et al., 2023; Zouhar et al., 2023). However, these emperically findings are in conflicted with other later studies Cognetta et al. (2024); Schmidt et al. (2024). On the theoretical side, Rajaraman et al. (2024) examined tokenization through the lens of unigram models, motivated by the observation made by Makkuva et al. (2024) that transformers struggles to learn 2nd-order Markov chains. We, however, do not observe this phenomenon in our experiment. As such, our work on bias due to tokenization is not affected by their observation.
Tokenization and Perplexity. Our work relates to the statistical evaluation of LMs, where we provide an algorithm to directly evaluate the character-level perplexity $p(x^{N}_{n+1}|x^{n}_{1})$, using a tokenized LM. In terms of token-level perplexity evaluation, some recent studies (Cao & Rimell, 2021; Chirkova et al., 2023) have suggested using stochastic tokenization (Provilkov et al., 2019) at test time to evaluate perplexity scores of LMs ($p(t^{i}_{1})$). However, these evaluations were done on LMs trained with deterministic tokenization which could be suboptimal as demonstrated by our examination of undefined states in Section 2. As such, by utilizing our approach, one can obtain a much more accurate insights on LMs evaluation.
Related Algorithms. Our algorithm is inspired from the literature of universal compression such as prediction by partial matching (Cleary & Witten, 1984) and context-tree weighting (Willems et al., 1995), which have been applied for text prediction but for much simpler settings without any tokenization involved. Recently, Minixhofer et al. (2024); Liu et al. (2023a) propose tokenization adaptation methods, which still requires a heuristic optimization that complicates the training pipeline. Some recent studies have proposed method to target the problem of language models encountering difficulties generating text near prompt boundaries (Dagan et al., 2024; guidance ai, 2023), which bears some resemblance to our proposed algorithm. These methods, however, are heuristic and only applicable to certain scenarios. On the other hand, our bias removal algorithm is theoretically correct, versatile for various situations, and enables conversion between token-free and tokenized LMs due to its accurate representation of conditional sampling distributions.
Appendix B Supporting Theorems on Tokenization
This section provides supporting theorems for the proof of the main results. We first remind the readers that the set $\mathcal{S}(x^{n}_{1})$ corresponds to the set of all strings that contains $x^{n}_{1}$ as a prefix. Similarly, the event set $\mathcal{S}(t^{i}_{1})$ corresponds to the set of all strings whose first $i$ tokens are $t^{i}_{1}$. Consider when $t^{i}_{1}=\mathrm{encode}(x^{n}_{1})$, it should be noted that the two sets $S(t^{i}_{1})$ and $S(x^{n}_{1})$ are not guaranteed to be equivalent. That is because the subsequent characters after $x^{n}_{1}$ can affect the tokenization within the first $n$ character. We illustrate this in more details in the following example.
Example. Consider the Markov chain example in Section 2, where $\mathcal{V}=\{``AA",``A",``B"\}$. Then, the string $s_{1}=``AABAABAB"$, then $s_{1}\in\mathcal{S}(x_{1}=``A")$ and $s_{1}\in\mathcal{S}(t_{1}=``AA")$ since the first character of $s_{1}$ is ``$A"$ and the first token of $s_{1}$ is ``$AA"$. On the other hand, $s_{1}\notin\mathcal{S}(t_{1}=\mathrm{encode}(x_{1})=``A")$ since its first token is ``$AA"$, not ``$A"$.
We introduce the following proposition that contains two facts regarding the maximum prefix encoding process.
Proposition B.1.
Let $s$ be a string with the prefix $x^{n}_{1}$ ($x^{n}_{1}\in\mathrm{prefix}(s)$). Define the minimal superstring $r$ to be the prefix of $s$ with the fewest tokens that contains $x^{n}_{1}$ as a prefix: $r=\mathrm{argmin}_{r}(k|t^{k}_{1}=\mathrm{encode}(r)\wedge x^{n}_{1}\in\mathrm{prefix}(r)\wedge r\in\mathrm{prefix}(s))$. Then, we have the followings:
  1. 1.
    For $1\leq i<k$, $\mathrm{encode}(s)_{i}=\mathrm{encode}(x^{n}_{1})_{i}$. Furthermore, when $r=x^{n}_{1}$, we also have $\mathrm{encode}(s)_{k}=\mathrm{encode}(x^{n}_{1})_{k}$.
  2. 2.
    Let $\ell$ be the number of tokens in $\mathrm{encode}(x^{n}_{1})$, then we have $\mathrm{decode}(\mathrm{encode}(x^{n}_{1})^{\ell}_{k})\in\mathrm{prefix}(\mathrm{decode}(\mathrm{encode}(s)_{k}))$.
Proof.
(Result 1.) Proof by contradiction. Let $s$ be the counter-example with the fewest number of tokens. Assume that for $1\leq i<k$, $\mathrm{encode}(s)_{i}\neq\mathrm{encode}(x^{n}_{1})_{i}$. Let $j$ be the smallest of such $i$.
Consider $\mathrm{encode}(s)_{j}$ and $\mathrm{encode}(x^{n}_{1})_{j}$.
  • Case 1: $|\mathrm{decode}(\mathrm{encode}(s)^{j}_{1})|<|x^{n}_{1}|$.
    • Case 1.a: $|\mathrm{decode}(\mathrm{encode}(s)_{j})|<|\mathrm{decode}(\mathrm{encode}(x^{n}_{1})_{j})|$. This leads to a contradiction, since $x^{n}_{1}$ is a prefix of $s$, therefore a longest prefix matching algorithm would always generate the longer token ($\mathrm{encode}(x^{n}_{1})_{j}$) over the shorter one ($\mathrm{encode}(s)_{j}$) when it is available.
    • Case 1.b: $|\mathrm{decode}(\mathrm{encode}(s)_{j})|>|\mathrm{decode}(\mathrm{encode}(x^{n}_{1})_{j})|$. This leads to a contradiction, since $\mathrm{concat}(\mathrm{encode}(s)^{j}_{1})$ is a prefix of $x^{n}_{1}$ (Case 1 assumption), therefore a longest prefix matching algorithm would always generate the longer token ($\mathrm{encode}(s)_{j}$) over the shorter one ($\mathrm{encode}(x^{n}_{1})_{j}$) when it is available.
    • Case 1.c: $|\mathrm{decode}(\mathrm{encode}(s)_{j})|=|\mathrm{decode}(\mathrm{encode}(x^{n}_{1})_{j})|$. This means that the two tokens are the same, contradicting our initial assumption.
  • Case 2: $|\mathrm{decode}(\mathrm{encode}(s)^{j}_{1})|\geq|x^{n}_{1}|$. In this case, $r=\mathrm{decode}(\mathrm{encode}(s)^{j}_{1})$ is a superstring of $x^{n}_{1}$ implying that $k$ is at most $j$, which contradicts our initial assumption that $1\leq j<k$.
Finally, in the case $r=x^{n}_{1}$, this means $\mathrm{decode}(\mathrm{encode}(s)_{k})$ is a suffix of $x^{n}_{1}$. Since all the tokens before $k$ within $x^{n}_{1}$ has been matched, i.e. $\mathrm{encode}(s)_{i}=\mathrm{encode}(x^{n}_{1})_{i}$ for $1\leq i<k$, the last token must also match as the result (else, $|r|\neq|x^{n}_{1}|$, leads to contradiction), we have $\mathrm{encode}(s)_{k}=\mathrm{encode}(x^{n}_{1})_{k}$.
(Result 2.) The proof idea is that since $r$ contains $x^{n}_{1}$ and any tokens within $r$ and $x^{n}_{1}$ has been matched up to $k{-}1$, then what is left in $x^{n}_{1}$ must be in the last token in $r$ (which is the $k$th token of $r$). Formally, following Result 1, we have $\mathrm{decode}(\mathrm{encode}(x^{n}_{1})^{k-1}_{1})=\mathrm{decode}(\mathrm{encode}(s)^{k-1}_{1})$. Since $r$ has $k$ tokens in total and $x^{n}_{1}\in\mathrm{prefix}(r)$, this means that $\mathrm{decode}(\mathrm{encode}(s)_{k})$ must cover the rest of $x^{n}_{1}$, i.e. $\mathrm{decode}(\mathrm{encode}(x^{n}_{1})^{\ell}_{k})$. As the result, we must have $\mathrm{decode}(\mathrm{encode}(x^{n}_{1})^{\ell}_{k})\in\mathrm{prefix}(\mathrm{decode}(\mathrm{encode}(s)_{k}))$. ∎
We remind the reader the definition of invalid encoding below.
Definition B.2.
(Invalid Encodings) The list of tokens (an encoding) $t^{k}_{1}$ is invalid if $\mathrm{encode}(\mathrm{decode}(t^{k}_{1})){\neq}t^{k}_{1}$. Otherwise, it is a valid encoding.
Corollary B.3.
$\mathcal{S}(t^{k}_{1})=\emptyset$ if and only if $t^{k}_{1}$ is invalid.
Proof.
We prove each direction as follow.
  • If $\mathcal{S}(t^{k}_{1})=\emptyset$ then $t^{k}_{1}$ is invalid: Since $\mathcal{S}(t^{k}_{1})=\emptyset$, we know that there exist no string $s$ such that $\mathrm{encode}(s)^{k}_{1}=t^{k}_{1}$. As such, for $s=\mathrm{decode}(t^{k}_{1})$, we do not have $\mathrm{encode}(\mathrm{decode}(t^{k}_{1}))=t^{k}_{1}$, which proves the result.
  • If $t^{k}_{1}$ is invalid then $\mathcal{S}(t^{k}_{1})=\emptyset$: We prove this by showing that it is impossible to have a string with prefix $\mathrm{decode}(t^{k}_{1})$ whose first $k$ tokens are $t^{k}_{1}$. Let $\bar{s}=\mathrm{decode}(t^{k}_{1})$, we have any string $s\in\mathcal{S}(t^{k}_{1})$ must have $\bar{s}\in\mathrm{prefix}(s)$. Assume such string $s$ exists, we have $\mathrm{encode}(s)^{k}_{1}=t^{k}_{1}$. Let $x^{n}_{1}=\mathrm{decode}(t^{k}_{1})$, according to Proposition B.1, we have $r=\bar{s}$ and as the result, for $1\leq i\leq k$, we have $\mathrm{encode}(s)_{i}=t_{i}=\mathrm{encode}(x^{n}_{1})_{i}$, which contradicts our initial assumption. This completes the proof.
Appendix C Proof of Proposition 2.3 in the Main Paper
Proposition 2.3 (Token-Induced Zero Probability) Let $t^{i}_{1}$ be a sequence of input tokens. For any invalid encoding $t^{i}_{1}$, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ and the conditional probability $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ is undefined. In the case $t^{i}_{1}$ is valid, then $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1}){=}0.0$ if $t^{i+1}_{1}$ is invalid. Furthermore, let $x^{n}_{1}{=}\mathrm{decode}(t^{i}_{1})$, then for any string $x^{N}_{n+1}$ such that $\mathrm{encode}(\mathrm{concat}(\mathrm{decode}(t^{i}_{1}),x^{N}_{n+1})){\neq}t^{i}_{1}$, then $P_{\mathrm{gt}}(x^{N}_{n+1}|t^{i}_{1}){=}0.0.$
Proof.
For the first two statements, we have:
  • For an invalid $t^{i}_{1}$ where $t^{i}_{1}\neq\mathrm{encode}(\mathrm{decode}(t^{i}_{1}))$, we have $\mathcal{S}(t^{i}_{1})=\emptyset$, as implied by Corollary B.3. As such, we have $P_{\mathrm{gt}}(t^{i}_{1}){=}0.0$ which leads $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})$ to be an undefined conditional probability .
  • For a valid $t^{i}_{1}$ but invalid $t^{i+1}_{1}$, we know that $P_{\mathrm{gt}}(t^{i+1}_{1}){=}0.0$, which results in $P_{\mathrm{gt}}(t_{i+1}|t^{i}_{1})=0.0$.
For the last statement, we first note the following:
  1. 1.
    Let $s^{\prime}=\mathrm{decode}(t^{i}_{1})$, note that $P_{\mathrm{gt}}(x^{N}_{n+1},t^{i}_{1})=P_{\mathrm{gt}}(s^{\prime},x^{N}_{n+1},t^{i}_{1})=P_{\mathrm{gt}}(\bar{s},t^{i}_{1})$ where $\bar{s}=\mathrm{concat}(s^{\prime},x^{N}_{n+1})=x^{N}_{1}$.
  2. 2.
    Consider $P_{\mathrm{gt}}(\bar{s},t^{i}_{1})=P_{\mathrm{gt}}(\bar{s})P_{\mathrm{gt}}(t^{i}_{1}|\bar{s})$, we will prove that $P_{\mathrm{gt}}(t^{i}_{1}|\bar{s})=0.0$ if $\mathrm{encode}(\bar{s})^{i}_{1}\neq t^{i}_{1}$.
The idea is that, when we append additional characters to $\bar{s}$, it will never change the tokenization such that the first $i$ tokens are $t^{i}_{1}$. Formally:
  • Let $j$ be the first position such that $\mathrm{encode}(\bar{s})_{j}\neq t_{j}$ then we know that $|\mathrm{decode}(\mathrm{encode}(\bar{s})_{j}))|>|\mathrm{decode}(t_{j})|$ (Proposition B.1 (Result 2)).
  • Following Proposition B.1 (Result 2), let $s^{*}\in\mathcal{S}(\bar{s})$, then we know that $\mathrm{decode}(\mathrm{encode}(\bar{s})_{j})$ must be a substring of within another longer token (it cannot be broken down). This completes the proof.
Finally, we note that $P_{\mathrm{gt}}(t^{i}_{1})=0.0$ does not implies $\mathrm{encode}(\mathrm{decode}(t^{i}_{1}))\in\mathcal{V}$, since it can be due to the original distribution on the character domain. A classic example for this is a Markov model with an absorption state. ∎
Appendix D Proof of Proposition 3.1 and Corollary 3.2 in the Main Paper
Proposition 3.1 Let $s^{*}=x^{n}_{1}$, where $t^{i}_{1}=\mathrm{encode}(s^{*})=\mathrm{encode}(x^{n}_{1})$. Then we have $\mathcal{S}(t^{i}_{1})\subset\mathcal{S}(x^{n}_{1})$, i.e. for any string $s$ where $t^{i}_{1}=\mathrm{encode}(s)^{i}_{1}$, we have $P(x^{n}_{1}|t^{i}_{1})=1.0$. In the case $t_{i}\in\mathcal{V}^{*}$, then we also have that $\mathcal{S}(t^{i}_{1})=\mathcal{S}(x^{n}_{1})$, i.e. any string $s$ where $x^{n}_{1}\in\mathrm{prefix}(s)$ must have the first $i$ tokens as $t^{i}_{1}$ and $P(t^{i}_{1}|x^{n}_{1})=1.0$.
Proof.
We prove each case as follow.
1) General Case: Since there exists a string $s$ with prefix $x^{n}_{1}$ where $\mathrm{encode}(s)^{i}_{1}\neq t^{i}_{1}$, which follows directly from our 1st order Markov chain example in the main paper, i.e. the string $s=``AA"$ has ``$A"$ as prefix but have the $t_{1}=``AA"\neq``A"$. On the other hand, we know that any string $s$ that has the first $i$ tokens as $t^{i}_{1}$ must have the first $n$ characters as $x^{n}_{1}$, as such $\mathcal{S}(t^{i}_{1})\subset\mathcal{S}(x^{n}_{1})$ and $P(x^{n}_{1}|t^{i}_{1})=1.0$.
2) $t_{i}\in\mathcal{V}^{*}$: The proof idea is that, since $t_{i}$ cannot be a part of any token in $\mathcal{V}$, it is impossible to merge it by appending additional characters after $t_{i}$. Formally, similar to Proposition B.1:
  • For any string $s\in\mathcal{S}(\mathrm{decode}(t^{i}_{1}))$, let $\ell$ be the number of tokens in the minimal superstring $r$ of $s$ that contains $x^{n}_{1}$ as a prefix.
  • Following Proposition B.1 (Result 2), we know that $t_{i}$ must be a substring of $\mathrm{decode}(\mathrm{encode}(s)_{\ell})$.
  • Due to $t_{i}\in\mathcal{V}^{*}$, then $t_{i}=\mathrm{encode}(s)_{\ell}$. We also know from Proposition B.1 (Result 1) that $\mathrm{encode}(s)_{i}=t_{i}$ for $1\leq i<\ell$, this means that $\ell=i$. This gives us $t^{i}_{1}=\mathrm{encode}(s)^{i}_{1}$ and $P(t^{i}_{1}|x^{n}_{1})=1.0$.
This completes the proof. ∎
Remarks. We briefly note that the condition $t_{i}\in\mathcal{V}^{*}$ is the sufficient condition. In general, any token sequence $t^{i}_{1}$ that satisfies $\mathcal{S}(t^{i}_{1})=\mathcal{S}(x^{n}_{1})$ will have $P(t^{i}_{1}|x^{n}_{1})=1.0$. Developing an efficient algorithm to verify whether $\mathcal{S}(t^{i}_{1})=\mathcal{S}(x^{n}_{1})$ remains a task for future research.
Corollary 3.2 Following Proposition 3.1, suppose $t_{i}\in\mathcal{V}^{*}$ then we have $P(x^{N}_{n+1}|x^{n}_{1}{)}{=}P(x^{N}_{n{+}1}|t^{i}_{1})$. Similarly, we also have $P(t^{j}_{i+1}|x^{n}_{1}{)}{=}P(t^{j}_{i+1}|t^{i}_{1})$.
Proof.
For the first case, we prove through the following equations:
$P(x^{N}_{n{+}1}|t^{i}_{1})$ $=P(x^{N}_{n{+}1}|t^{i}_{1},x^{n}_{1})$ (4)
$=\frac{P(x^{N}_{n+1},t^{i}_{1}|x^{n}_{1})}{P(t^{i}_{1}|x^{n}_{1})}$ (5)
$=\frac{P(x^{N}_{n+1}|x^{n}_{1})P(t^{i}_{1}|x^{n}_{1},x^{N}_{n+1})}{P(t^{i}_{1}|x^{n}_{1})}$ (6)
$=P(x^{N}_{n+1}|x^{n}_{1})$ (7)
where the first equality is due to $P(x^{n}_{1}|t^{i}_{1})=1.0$ and the last equality is due to $P(t^{i}_{1}|x^{n}_{1})=1.0$ for $t_{i}\in\mathcal{V}^{*}$.
Similarly, for the second case, we have:
$P(t^{j}_{i+1}|t^{i}_{1}{)}$ $=P(t^{j}_{i+1}|x^{n}_{1},t^{i}_{1})$ (8)
$=\frac{P(t^{j}_{i+1}|x^{n}_{1})P(t^{i}_{1}|x^{n}_{1},t^{j}_{i+1})}{P(t^{i}_{1}|x^{n}_{1})}$ (9)
$=P(t^{j}_{i+1}|x^{n}_{1}),$ (10)
which completes the proof. ∎
Appendix E Proof for The Bias Removal Method
E.1 Refactoring
Our goal is to express the quantity $P(x^{N}_{n+1}|x^{n}_{1})$ using the tokenized LM that outputs the conditional probability $P(t_{i}|t^{i-1}_{1})$. Let $x^{n_{k}}_{1}\in\mathrm{prefix}(x^{n}_{1})$ where $t^{k}_{1}=\mathrm{encode}(x^{n_{k}}_{1})$ and $t_{k}\in\mathcal{V}^{*}$. Following Proposition 3.1, any string $s$ with prefix $x^{n_{k}}_{1}$ must have the first $k$ tokens as $t^{k}_{1}$. We now perform the following factorization:
$P(x^{N}_{n+1}|x^{n}_{1})$ $=P(x^{N}_{n+1}|x^{n_{k}}_{1},x^{n}_{n_{k}+1})$ (11)
$=\frac{P(x^{N}_{n+1},x^{n}_{n_{k}+1}|x^{n_{k}}_{1})}{P(x^{n}_{n_{k}+1}|x^{n_{k}}_{1})}$ (12)
$=\frac{P(x^{N}_{n_{k}+1}|x^{n_{k}}_{1})}{P(x^{n}_{n_{k}+1}|x^{n_{k}}_{1})}$ (13)
$=\frac{P(x^{N}_{n_{k}+1}|t^{k}_{1})}{P(x^{n}_{n_{k}+1}|t^{k}_{1})},$ (14)
where the last inequality is due to Corollary 3.2. Finally, we will use the Branch and Pass Algorithm to compute each term in (14) individually. Note that the algorithm does not require $t_{k}\in\mathcal{V}^{*}$. Here, we explicitly highlight the importance of having $t_{k}\in\mathcal{V}^{*}$, as it bridges between the character and token domain through Equation (14).
E.2 The Branch and Pass Algorithm
Overview. The Branch and Pass (BPTree) algorithm has two inputs: the query string value $x^{N}_{n_{k}+1}$ and the conditioning tokens $t^{k}_{1}$, which outputs the value $P(x^{N}_{n_{k}+1}|t^{k}_{1})$. Note that we do not require $t_{k}\in\mathcal{V}^{*}$ in the BPTree algorithm. First, using marginalization, we have the following:
$P(x^{N}_{n_{k}+1}|t^{k}_{1})$ $=\sum_{t\in\mathcal{V}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})$ (15)
$=\underbrace{\sum_{t{\in}\mathcal{T}_{\mathrm{b_{val}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})}_{\mathrm{b_{val}}}+\underbrace{\sum_{t{\in}\mathcal{T}_{\mathrm{p_{val}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})}_{\mathrm{p_{val}}}$ (16)
where $\mathcal{T}_{\mathrm{b_{val}}}{=}\{t|t{\in}\mathcal{V}{\wedge}x^{N}_{n_{k}+1}{\in}\mathrm{prefix}(\mathrm{decode}(t))\}$ and $\mathcal{T}_{\mathrm{p_{val}}}{=}\{t|t\in\mathcal{V}\wedge x^{N}_{n_{k}+1}{\notin}\mathrm{prefix}(\mathrm{decode}(t))\}$, where $\mathcal{T}_{\mathrm{b_{val}}}\cap\mathcal{T}_{\mathrm{p_{val}}}=\emptyset$, are the corresponding set we iterates over each sum in the branch step, i.e. $\mathrm{b_{val}}$, and pass step, i.e. $\mathrm{p_{val}}$, respectively. Intuitvely, $\mathcal{T}_{\mathrm{b_{val}}}$ is the set of tokens that have a prefix $x^{N}_{n_{k}+1}$ and $\mathcal{T}_{\mathrm{p_{val}}}$ is the ones that do not. We now show how to compute each term separately using a tokenized LM that outputs $P(t_{i+1}|t^{i}_{1})$.
Branch Step. Note that $\mathrm{b_{val}}$ can be described as the probability that, given the list of previous tokens $t^{k}_{1}$, the next token of the string $s$ has $x^{N}_{n_{k}+1}$ as a prefix. As such, to obtain $\mathrm{b_{val}}$, we first compute $P(t_{k+1}=t|t^{k}_{1})$ for all $t\in\mathcal{V}$ using one model run, then sum the probabilities corresponds to all tokens whose prefix is $x^{N}_{n_{k}+1}$. Specifically:
$\mathrm{b_{val}}=\sum_{t\in\mathcal{T}_{\mathrm{b_{val}}}}P(t_{k+1}=t|t^{k}_{1}),$ (17)
Proof.
To see this, for each summand of $\mathrm{b_{val}}$ in Eq.(16), we have:
$P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})$ $=P(t_{k+1}=t|t^{k}_{1})\times P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}{=}t)$ (18)
$=P(t_{k+1}=t|t^{k}_{1}),$ (19)
where the second equation is due to $x^{N}_{n_{k}+1}{\in}\mathrm{prefix}(t)$ so $P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}=t){=}1.0$. This concludes the proof. ∎
Pass Step. In contrast to $\mathrm{b_{val}}$, the $\mathrm{p_{val}}$ is the probability that, given the list of previous (valid) tokens $t^{k}_{1}$, the string $s$ has the subsequent string as $x^{N}_{n_{k}+1}$ that is not a prefix of the next token. We will show that under the maximum prefix encoding, we compute the value $\mathrm{p_{val}}$ as follow:
$\mathrm{p_{val}}=P(t_{k+1}=t|t^{k}_{1})\times P(x^{N}_{n_{k+1}+1}|t^{k}_{1},t_{k+1}=t),$ (20)
where $t=\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$ and $x^{n_{k+1}}_{n_{k}+1}=\mathrm{decode}(t)$. That is, during the passing step, there are two subroutines:
  1. 1.
    Obtain the next token $t$ within $x^{N}_{n_{k}+1}$ and compute $P(t_{k+1}=t|t^{k}_{1})$. If $x^{N}_{n_{k}+1}=\mathrm{decode}(t)$, then returns $0.0$ since this is not allowed according to the condition required in $\mathcal{T}_{\mathrm{p_{val}}}$.
  2. 2.
    Else, recursively compute $P(x^{N}_{n_{k+1}+1}|t^{k}_{1},t_{k+1}=t)$.
Proof.
We first rewrite $\mathcal{T}_{\mathrm{p_{val}}}$ as the union of the following disjoint sets, i.e $\mathcal{T}_{\mathrm{p_{val}}}{=}\mathcal{T}_{\mathrm{p_{val,1}}}{\cup}\mathcal{T}_{\mathrm{p_{val,2}}}$ and $\mathcal{T}_{\mathrm{p_{val,1}}}{\cap}\mathcal{T}_{\mathrm{p_{val,2}}}{=}\emptyset$, where:
  • $\mathcal{T}_{\mathrm{p_{val,1}}}=\{t|t\in\mathcal{V}\wedge x^{N}_{n_{k}+1}\notin\mathrm{prefix}(\mathrm{decode}(t))\wedge\mathrm{decode}(t)\notin\mathrm{prefix}(x^{N}_{n_{k}+1})\}$: the subset of $\mathcal{T}_{\mathrm{p_{val}}}$ where the string represented by the token $t$ is not a prefix of $x^{N}_{n_{k}+1}$.
  • $\mathcal{T}_{\mathrm{p_{val,2}}}=\{t|t\in\mathcal{V}\wedge x^{N}_{n_{k}+1}\notin\mathrm{prefix}(\mathrm{decode}(t))\wedge\mathrm{decode}(t)\in\mathrm{prefix}(x^{N}_{n_{k}+1})\}$: the subset of $\mathcal{T}_{\mathrm{p_{val}}}$ where string represented by the token $t$ is actually a prefix of $x^{N}_{n_{k}+1}$.
As such, we have:
$\mathrm{p_{val}}=\sum_{t\in\mathcal{T}_{\mathrm{p_{val}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=\underbrace{\sum_{t\in\mathcal{T}_{\mathrm{p_{val,1}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})}_{0.0}+\sum_{t\in\mathcal{T}_{\mathrm{p_{val,2}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1}),$ (21)
where we note that for $t\in\mathcal{T}_{\mathrm{p_{val,1}}}$, we have $P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}=t)=0.0$ since the conditional $t_{k+1}=t$ implies that the string represented by $t$ must be a prefix of $x^{N}_{n_{k}+1}$, which leads to the first sum being $0.0$.
For the second term, i.e. $t\in\mathcal{T}_{\mathrm{p_{val,2}}}$, if $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, by following Proposition 2.3, we have $P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=0.0$ since any tokenization that is not a valid maximum prefix encoding must happen with probability $0.0$. This stands corrected as long as the output probability of our LM follows Proposition 2.3 (see Section G.2).
As a result, we are left with the token $t=\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, which is the Equation (20) after applying chain rule of probability. ∎
Base Case. We note that the base case of our algorithm corresponds to the situation where $x^{N}_{n_{k}+1}=\mathrm{decode}(t)$. In this scenario, we only needs to compute $\mathrm{b_{val}}$ (branching step) while $\mathrm{p_{val}}=0.0$.
Complexity Analysis. The complexity of our algorithm (number of inferences on the language model) scales with the length of the the query string, i.e. $N-n_{k}$. Note that the complexity of the summation at the Branching step is relatively cheap compared to the runtime of the language model.
Appendix F Converting Token-Free Language Model to Tokenized Language Model.
We introduce an algorithm to compute $P(t_{k+1}|t^{k}_{1})$ using a token-free language model $P(x^{N}_{n+1}|x^{n}_{1})$, despite having no access to any tokenized LM. This approach enables theoretical conversion of a token-free model to a tokenized one. The method involves two stages. First, we refactor the conditional probability similar to the technique presented in Section E. Next, we aggregate the probabilities of all possible strings leading to the desired tokenization. It is important to note that a Markov chain is a special type of autoregressive model, meaning this method can be employed to effortlessly calculate Markov chain transition matrices within the tokenized domain.
F.1 Refactoring
Consider the probability $P(t_{i+1}|t^{i}_{1})$ that we would like to expressed using $P(x^{N}_{n+1}|x^{n}_{1})$. Let $t_{k}$ be the last token within $t^{i}_{1}$ such that $t_{k}\in\mathcal{V}^{*}$. We now perform the following factorization:
$P(t_{i+1}|t^{i}_{1})$ $=\frac{P(t^{i+1}_{k+1}|t^{k}_{1})}{P(t^{i}_{k+1}|t^{k}_{1})}$ (22)
$=\frac{P(t^{i+1}_{k+1}|x^{n_{k}}_{1})}{P(t^{i}_{k+1}|x^{n_{k}}_{1})},$ (23)
where $x^{n_{k}}_{1}=\mathrm{decode}(t^{k}_{1})$. The second equality is due to Corollary 3.2. Each term can then be computed using the aggregation procedure shown next.
F.2 Aggregation.
In this step, we would like to compute $P(t^{i}_{k+1}|x^{n_{k}}_{1})$ where $\mathrm{encode}(x^{n_{k}}_{1})=t^{k}_{1}$ and $t_{k}\in\mathcal{V}^{*}$, using the token-free representation $P(x_{n+1}|x^{n}_{1})$. Here, we denote $\mathrm{decode}(t^{i}_{k+1})=x^{n_{i}}_{n_{k}+1}$ and $M=\max_{t\in\mathcal{V}}|\mathrm{decode}(t)|$ be the length of the longest token in $V$ and $\Omega=\mathcal{A}^{M}$ is the enumeration of all string of length $M$.
Computing $P(t^{i}_{k+1}|x^{n_{k}}_{1})$ involves considering all possible strings $s$ with prefix $x^{n_{i}}_{1}$ and $t^{i}_{k+1}=\mathrm{encode}(s)^{i}_{k+1}$. Although iterating through every possible string is infeasible, we can restrict our search by only examining strings with length $|s|=n_{i}+M$, as any additional string beyond this point will not impact the tokenization of prefix $x^{n_{i}}_{1}$due to $M$ being the maximum token length. Formally, we will show that one can express $P(t^{i}_{k+1}|x^{n_{k}}_{1})$ as follows:
$P(t^{i}_{k+1}|x^{n_{k}}_{1}){=}\sum_{s^{\prime}\in\mathcal{A}^{M}}P(x^{n_{i}+M}_{n_{k}+1}{=}c_{1}(s^{\prime})|x^{n_{k}}_{1})\mathds{1}(t^{i}_{k+1}{=}\mathrm{encode}(c_{2}(s^{\prime}))^{i}_{k+1}),$ (24)
where $c_{1}(s^{\prime}):=\mathrm{concat}(x^{n_{i}}_{n_{k}+1},s^{\prime})$ and $c_{2}(s^{\prime}):=\mathrm{concat}(x^{n_{i}}_{1},s^{\prime})$. The first term can be computed using the given token-free LM, i.e. $P(x^{n_{i}+M}_{n_{k}+1}|x^{n_{k}}_{1})$. The second term is an indicator function that checks whether $t^{i}_{k+1}=\mathrm{encode}(s)^{i}_{k+1}$ and can be computed deterministically.
Proof.
We have:
$P(t^{i}_{k+1}|x^{n_{k}}_{1})$ $=P(t^{i}_{k+1},x^{n_{i}}_{n_{k}+1}|x^{n_{k}}_{1})$ (25)
$=\sum_{s^{\prime}\in\mathcal{A}^{M}}P(t^{i}_{k+1},x^{n_{i}+M}_{n_{k}+1}{=}c_{1}(s^{\prime})|x^{n_{k}}_{1})$ (26)
$=\sum_{s^{\prime}\in\mathcal{A}^{M}}P(x^{n_{i}+M}_{n_{k}+1}{=}c_{1}(s^{\prime})|x^{n_{k}}_{1})P(t^{i}_{k+1}|x^{n_{i}+M}_{1}=c_{2}(s^{\prime}))$ (27)
$=\sum_{s^{\prime}\in\mathcal{A}^{M}}P(x^{n_{i}+M}_{n_{k}+1}{=}c_{1}(s^{\prime})|x^{n_{k}}_{1})\mathds{1}(t^{i}_{k+1}{=}\mathrm{encode}(c_{2}(s^{\prime}))^{i}_{k+1})$ (28)
The rest is to prove the following equality:
$P(t^{i}_{k+1}|x^{n_{i}+M}_{1}=c_{2}(s^{\prime}))=\mathds{1}(t^{i}_{k+1}{=}\mathrm{encode}(c_{2}(s^{\prime}))^{i}_{k+1})$ (29)
We first note that the first $k$ tokens must be $t^{k}_{1}=\mathrm{encode}(x^{n_{k}}_{1})$ due to our condition that $t_{k}\in\mathcal{V}^{*}$. Since $M$ is the length of the longest token in $\mathcal{V}$, appending extra characters cannot change the tokenization happened for $x^{n_{i}}_{1}$. In other words, any string $s$ with prefix $c_{2}(s^{\prime})$ must have the same minimal superstring $r$ containing $x^{n_{i}}_{1}$ (see Proposition B.1). We then apply this principle to the two cases:
  • $t^{i}_{k+1}{=}\mathrm{encode}(c_{2}(s^{\prime}))^{i}_{k+1}$: In this case, we know that the string must contains the first $i$ tokens as $t^{i}_{1}$, hence $P(t^{i}_{k+1}|x^{n_{i}+M}_{1}=c_{2}(s^{\prime}))=1.0$
  • $t^{i}_{k+1}{\neq}\mathrm{encode}(c_{2}(s^{\prime}))^{i}_{k+1}$: In contrast, this case is equivalent to $P(t^{i}_{k+1}|x^{n_{i}+M}_{1}=c_{2}(s^{\prime}))=0.0$ since we are sure that the string do not contains the tokens $t^{i}_{k+1}$.
This concludes the proof.
F.3 The Markov Chain Example.
We provide a detail computation of the Markov chain example in the main paper. Recall that in the original chain (in the character domain), we have the following:
$P(x_{2}=``A"|x_{1}=``A")$ $=\alpha$ (30)
$P(x_{2}=``B"|x_{1}=``A")$ $=1-\alpha$ (31)
$P(x_{2}=``A"|x_{1}=``B")$ $=\beta$ (32)
$P(x_{2}=``B"|x_{1}=``B")$ $=1-\beta$ (33)
We also assume the initial probability $\pi=\{\gamma,1-\gamma\}$ for ``$A"$ and ``$B"$ respectively. In the token domain, let first compute $P(t_{2}=``A"|t_{1}=``AA")$, where we do not have to do the refactoring step since we know that $t_{1}\in\mathcal{V}^{*}$. Following the Aggregation step, we have:
$P(t_{2}=``A"|t_{1}=``AA")$ $=P(x^{6}_{3}=``ABA"|x^{2}_{1}=``AA")+P(x^{6}_{3}=``ABB"|x^{2}_{1}=``AA")$ (34)
$=P(x^{5}_{3}=``AB"|x^{2}_{1}=``AA")$ (35)
$=\alpha(1-\alpha),$ (36)
where in the first equality, we do not include the case $x^{6}_{3}="AAA"$ and $x^{6}_{3}="AAB"$ since $\mathrm{encode}("AAA")_{1}=``AA"$ and $\mathrm{encode}("AAB")_{1}=``AA"$, which are not the token ``$A"$ that we are interested in. For other tokens and when $t_{1}=``B"$, the computation follows the same arguments.
We now consider the case $P(t_{2}=``B"|t_{1}=``A")$, we can refactor it as:
$P(t_{2}=``B"|t_{1}=``A")=\frac{P(t_{2}=``B",t_{1}=``A")}{P(t_{1}=``A")}$ (37)
We first compute $P(t_{1}=``A")$ using the aggregation step:
$P(t_{1}=``A")$ $=P(x^{3}_{1}=``ABB")+P(x^{3}_{1}=``ABA")$ (38)
$=P(x^{2}_{1}=``AB")$ (39)
$=\gamma(1-\alpha),$ (40)
where we do again include the case $x^{6}_{3}="AAA"$ and $x^{6}_{3}="AAB"$ for the same reason above. For $P(t_{2}=``A",t_{1}=``A")$ we have:
$P(t_{2}{=}``B",t_{1}{=}``A")$ ${=}P(x^{4}_{1}{=}``ABAA")+P(x^{4}_{1}{=}``ABAB")+P(x^{4}_{1}{=}``ABBA")+P(x^{4}_{1}{=}``ABBB")$ (41)
${=}P(x^{2}_{1}=``AB")$ (42)
${=}\gamma(1-\alpha)$ (43)
which gives us $P(t_{2}=``B"|t_{1}=``A")=1.0$. Finally, in this specific case, since order of the Markov chain in the character domain is $1$, we do not need to consider the higher order of the Markov chain in the token domain.
Appendix G On Predictive Distribution of LMs
In practice, our LM does not follow Proposition 2.3 due to softmax activations. As such, in our BPTree algorithm, when $t\in\mathcal{T}_{\mathrm{p_{val,2}}}$ and $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, then $P_{\theta}(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})$ may not be $0.0$ (where $\theta$ is the model weights). Eventually, this can potentially increase the complexity of our BPTree algorithm during the Passing step.
In the first part of this section, we show that given any tokenized LM, we can force its output probabilites to obey Proposition 2.3, without any loss in terms of perplexity score on the token domain. In the second part, for completeness, we will show that having a tokenized LM satifying Propostion 2.3 will guarantee the correctness of the Passing step in our BPTree algorithm.
Finally, before going to the method, we remind the readers that Proposition 3.1 and Corollary 3.2 are factually correct and hold for all $\theta$. As such, the refactoring step holds regardless.
G.1 Truncate-Renormalization Process
We justify the assumption that our tokenized language model $P_{\theta}(t_{i+1}|t^{i}_{1})$ follows Proposition 2.3. The idea is that we can turn a language model that does not follow Proposition 2.3 to the one that does while guaranteeing that the new model will always results in a higher token-level perplexity score.
We first introduce Proposition G.1. In this proposition, we are given a target discrete probability distribution $p$ where we know some of the values will not happen, says $\Phi^{*}$. Assume that we have another distribution $q$ that approximates $p$, then we can produce another distribution $q^{*}$ that is closer to $p$ in terms of KL divergence by setting corresponding probabilities of $q$ in $\Phi^{*}$ to $0.0$ and renormalize it (similar to rejection sampling).
Proposition G.1.
Given a discrete distribution $p=\{p_{1},p_{2},...,p_{m}\}$ and $q=\{q_{1},q_{2},...,q_{m}\}$ with $q_{i}>0.0$ for all $i$. Let $\Phi=\{i\in\mathbb{Z}|p_{i}=0.0\}$ and $\Phi^{*}\subseteq\Phi$, we define $q^{*}=\{q^{*}_{1},q^{*}_{2},...,q^{*}_{m}\}$ where $q^{*}_{i}=0.0$ for $i\in\Phi^{*}$, and $q^{*}_{j}=q_{j}/(\sum_{l\notin\Phi^{*}}q_{l})$. Then we have:
$D_{\mathrm{KL}}(p||q^{*})\leq D_{\mathrm{KL}}(p||q),$ (44)
which implies that $q^{*}$ is closer to $p$ than $q$. We refer the process of producing $q^{*}$ as truncate-renormalization (TR).
Proof.
Let $Z=(\sum_{l\notin\Phi}q_{l})$ is the normalizing factor in $q^{*}$. Note that $Z\leq 1$ and as such $\log(Z)\leq 0$. The proof is as follows:
$D_{\mathrm{KL}}(p||q^{*})$ $=\sum_{i}p_{i}\log\left(\frac{p_{i}}{q^{*}_{i}}\right)$ (45)
$=\sum_{i\notin\Phi^{*}}p_{i}\log\left(\frac{p_{i}}{q^{*}_{i}}\right)\quad\text{, use }0\log 0=0.0$ (46)
$=\sum_{i\notin\Phi^{*}}p_{i}\log\left(\frac{p_{i}}{q_{i}/Z}\right)$ (47)
$=\left[\sum_{i\notin\Phi^{*}}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)\right]+\log(Z)$ (48)
$\leq\sum_{i\notin\Phi^{*}}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)=D_{\mathrm{KL}}(p||q),$ (49)
which completes the proof. ∎
Applying this to our scenario, where we are given any autoregressive language models $\hat{P}_{\theta}(t_{i+1}|t^{i}_{1})$ that does not follow Proposition 2.3 (due to the softmax activations), we can perform the TR process (since we know which encoding is invalid) to obtain a new distribution ${P}_{\theta}(t_{i+1}|t^{i}_{1})$, which is guaranteed to be closer to the ground-truth model ${P_{\mathrm{gt}}}(t_{i+1}|t^{i}_{1})$. As the results, we are guaranteed that the token-level perplexity score of ${P}_{\theta}(t_{i+1}|t^{i}_{1})$ is always lower than or equal to $\hat{P}_{\theta}(t_{i+1}|t^{i}_{1})$.
G.2 On Passing Step in Branch and Pass Algorithm.
Once our tokenized LM follows Proposition 2.3, intuitively it can be regarded as an autogressive source (in the token domain) so it does not alternate the correctness of the Passing step. This section provides a proof for completeness.
Note that we only need to consider the second term of $\mathrm{p_{val}}$, i.e.:
$\mathrm{p_{val}}=\underbrace{\sum_{t\in\mathcal{T}_{\mathrm{p_{val,1}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})}_{0.0}+\sum_{t\in\mathcal{T}_{\mathrm{p_{val,2}}}}P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1}),$ (50)
as the other term is always $0.0$ regardless. Specifically, we will prove that, during the Pass step, when $t\in\mathcal{T}_{\mathrm{p_{val,2}}}$ and $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$, then $P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=0.0$ given that our LM follows Proposition 2.3. If $t^{k+1}_{1}$ is an invalid encoding, we immediately have $P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})=0.0$, so we only consider when $t^{k+1}_{1}$ is a valid encoding. We first prove the following Proposition.
Proposition G.2.
If $t^{k}_{1}$ is invalid, then $t^{k+1}_{1}$ is also invalid.
Proof.
Proof by contradiction, suppose $t^{k+1}_{1}$ is valid then $\mathcal{S}(t^{k+1}_{1})\neq\emptyset$. However, any string $s\in\mathcal{S}(t^{k+1}_{1})$ must also be in $\mathcal{S}(t^{k}_{1})$. But $\mathcal{S}(t^{k+1}_{1})=\emptyset$ (Corollary B.3). As such, $t^{k+1}_{1}$ is invalid. ∎
Since $x^{N}_{n_{k}+1}$ has totally $m=N-(n_{k}+1)$ characters, it can have maximum $m$ tokens. Let $\mathcal{C}=\{\tau\in\mathcal{V}^{m-1}|x^{N}_{n_{k}+2}\in\mathrm{prefix}(\mathrm{decode}(\tau))\}$ be the set of all possible sequence of $m-1$ tokens (both valid and invalid). We then have:
$P(x^{N}_{n_{k}+1},t_{k+1}=t|t^{k}_{1})$ $=\sum_{\tau\in\mathcal{C}}P(x^{N}_{n_{k}+1},t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})$ (51)
$=\sum_{\tau\in\mathcal{C}}P(t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})P(x^{N}_{n_{k}+1}|t^{k}_{1},t_{k+1}=t,t^{m}_{k+2}=\tau)$ (52)
Then, it is sufficient to show that $P(t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})=0.0$ for any $\tau\in\mathcal{C}$ as long as $t$ does not follow the maximum prefix encoding, i.e. $t\neq\mathrm{encode}(x^{N}_{n_{k}+1})_{1}$. We have the followings:
  • Since $t_{k+1}=t$ does not follow the maximum prefix encoding, then $\mathrm{encode}(\mathrm{decode}(t^{k+m}_{1})){\neq}t^{k+m}_{1}$ since we can always change the tokenization at $t_{k+1}$.
  • Let $t_{k+j}$ be the first token where $t^{k+j}_{1}$ is invalid. Then follow Proposition G.2, since $t^{k+j}_{1}$ is invalid, then any extra tokens after $t^{k+j}_{1}$ must also be invalid.
As such, we have: $P(t_{k+j}|t^{k+j-1}_{1})=0.0$, which leads to $P(t_{k+1}=t,t^{k+m}_{k+2}=\tau|t^{k}_{1})=0.0$, which proves the results.